OpenStudy (anonymous):

TRUE OR FALSE: The following is an arithmetic sequence. 1/2 , 1 , 3/2 , 2 ,...

jimthompson5910 (jim_thompson5910):

Take each term, and subtract off the previous term So 2 - 3/2 = ??? 3/2 - 1 = ??? 1 - 1/2 = ???

jimthompson5910 (jim_thompson5910):

What do you get for each result?

OpenStudy (anonymous):

0.5 1 0.5 ?

jimthompson5910 (jim_thompson5910):

2 - 3/2 = 4/2 - 3/2 = (4-3)/2 = 1/2 So 2 - 3/2 = 1/2 --------------------------------------- 3/2 - 1 = 3/2 - 2/2 = (3-2)/2 = 1/2 So 3/2 - 1 = 1/2 --------------------------------------- 1 - 1/2 = 2/2 - 1/2 = (2-1)/2 = 1/2 which means 1 - 1/2 = 1/2

jimthompson5910 (jim_thompson5910):

Notice how each difference is 1/2, so this tells us that this is an arithmetic sequence.

OpenStudy (anonymous):

thank you! :)

jimthompson5910 (jim_thompson5910):

yw :)